OpenStudy (jmartinez638):

Solve the differential equation y'= 2x^3 + 4x - 5.

OpenStudy (irishboy123):

go for it !! and post here

OpenStudy (prettygirl_shynice):

ill help

OpenStudy (jmartinez638):

Yeah, idk really what to do.

OpenStudy (irishboy123):

go @Prettygirl_shynice !!

OpenStudy (jmartinez638):

lol

OpenStudy (jmartinez638):

Oh no. I understand what's happening, and it is difficult

OpenStudy (jmartinez638):

A normal seperable equation would have, for example, an x and a y, and a dx and a dy. SO, since this is y', we must find y, seperate the equations, and integrate both sides.

OpenStudy (jmartinez638):

Sound about right?

OpenStudy (jmartinez638):

what do you think @Prettygirl_shynice

OpenStudy (tkhunny):

Have you considered first solving the characteristic equation?

OpenStudy (jmartinez638):

Not really. Can you explain?

OpenStudy (tkhunny):

You should not have been given this problem without knowing what a characteristic equation is. At least SHOW the characteristic equation or we really cannot proceed.

OpenStudy (tkhunny):

Once you find the roots of the equation, you can simply write down general solutions in terms of exponentials and sine and cosine.

OpenStudy (solomonzelman):

couldn't you just integrate?

OpenStudy (jmartinez638):

\[\frac{ x(x^3+4x-10) }{ 2 } + c\]

OpenStudy (solomonzelman):

you got that when integrating??

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \int x^n=\frac{x^{n+1}}{n+1}\color{grey}{+C} }\)

OpenStudy (zarkon):

jmartinez638's answer is correct...though it is unconventionally written

OpenStudy (solomonzelman):

oh.... -:( :)

OpenStudy (solomonzelman):

wouldn't write it like this myselft, but ...

OpenStudy (zarkon):

\[\frac{ x(x^3+4x-10) }{ 2 } + c=\frac{x^4+4x^2-10x }{ 2 } + c=\frac{x^4}{2}+2x^2-5+c\]

OpenStudy (solomonzelman):

yes, yes...

OpenStudy (jmartinez638):

\[\frac{ x^4 }{ 2 } +2x^2 -5x + \]

OpenStudy (jmartinez638):

Sorry, plus a constant

OpenStudy (zarkon):

left off the x \[\frac{ x(x^3+4x-10) }{ 2 } + c=\frac{x^4+4x^2-10x }{ 2 } + c=\frac{x^4}{2}+2x^2-5x+c\]

OpenStudy (jmartinez638):

Yeah, I 'simplified' it when I should have left it

OpenStudy (jmartinez638):

What next? I really haven't come across this specific problem, so

OpenStudy (irishboy123):

incroyable @ganeshie8

ganeshie8 (ganeshie8):

Solving a differential equation means finding a function "y" that satisfies the given equation. y = x^4/2+2x^2-5x satisfies the given differential equation, so this is a solution. All the solutions are given by y = x^4/2+2x^2-5x + c

OpenStudy (tkhunny):

Whoops! What was I thinking?